Krishna travels from P to Q by bus and returns from Q to P by bike in 10 hours. If he travels both ways by bus he saves 2 hours. What is the time taken to cover both ways by bike?

Answer by macston(5194)

.
If he saves 2 hours going both ways by bus, it takes 8 hours to make the entire trip by bus, or 4 hours in each direction. If 4 hours of the ten hours was the bus portion, it took him 6 hours to bike in one direction, and would take 12 hours to cover both ways by bike.
Let x=time by bike for entire trip and y=time by bus for entire trip
.
(1/2)x+(1/2)y=10 hours
y=10 hours-2 hours=8 hours
.
(1/2)x+(1/2)(8 hours)=10 hours
(1/2)x+4hrs=10 hrs
(1/2)x=6hrs
x=12 hours